A planet has a mass of 6.16 × 1023 kg and a radius of 3.59 × 106 m. (a) What is the acceleration due to gravity on this planet? (b) How much would a 51.1-kg person weigh on this planet?
Answer #1

a)

g = GM m / R^2

g = 6.67* 10^11* 6.16* 10^23 / (3.59* 10^6)^2 = 3.188 m/s^2

=======

b)

weight = mg = 51.1* 3.188 = 162.907 N

=======

Comment before rate in case any doubt, will reply for sure.. goodluck